# LiveMCPBench: Can Agents Navigate an Ocean of MCP Tools?

### v1

@@ C.2 Annotation Principles
Appendix D Details of LiveMCPTool Collection
D.1 Toolset Statistics
Figure 8: Distribution of tools in LiveMCPTool, categorized into 8 distinct types (total: 527 tools).
Figure 9: Distribution of servers in LiveMCPTool, categorized into 8 distinct types (total: 70 servers).
The statistics of LiveMCPTool’s tools and servers are illustrated in Figures 8-9. The collected toolset is categorized into eight distinct classes:
  1. 1.
@@ D.1 Toolset Statistics
Appendix E Implementation Details
Figure 10: An illustrative case of evaluation failure in LiveMCPEval. The assessment was conducted using DeepSeek-V3 on the completion trajectory from Claude-Sonnet-4. The evaluator model erroneously concluded the task was successful based solely on the agent’s file creation action, while failing to recognize that the agent did not actually acquire the required information.
E.1 Computing Resources and Private Models
@@
  • OpenRouter: GPT-4.1, GPT-4.1-Mini, DeepSeek-R1-0528, DeepSeek-V3-0324, Qwen3-235B-A22B and Qwen3-32B.
@@
To address suboptimal greedy decoding in certain reasoning models, we implemented a uniform temperature parameter of 0.7 across all experiments. This configuration introduces controlled stochasticity while maintaining result reliability for long-horizon tasks, as we observed that sporadic randomness has negligible cumulative impact on aggregate performance.
E.2 Tool Retrieval Configuration
For consistent retrieval performance, we established a standardized framework using Qwen2.5-72B-Instruct for tool summarization and Qwen3-Embedding-0.6B for embedding generation. To control for potential variance in this component, we maintained identical retrieval module parameters across all experimental conditions. Given the inherent temporal variability in tool outputs, we conducted all experiments within a tightly controlled window (July 20–27, 2025) to minimize fluctuations attributable to temporal factors.
Figure 11: Comparison of key points in DeepSeek-V3 and human annotations: Similar content despite different ordering.
Appendix F Evaluation Analysis
F.1 Case Study: Error Evaluation Examples
@@
F.2 Human and LLM Key Points Examples
@@
Appendix G Case Study: Error Examples
Figures 12-15 present concrete examples of four distinct error types: Query Error, Retrieve Error, Tool Error, and Other Error.
@@ Appendix G Case Study: Error Examples
Appendix H Prompts
H.1 MCP Copilot Agent Prompt
@@ H.2 LiveMCPEval Prompt
Figure 12: An illustration of Query Error: Discrepancy between the agent-generated query and the task’s required competencies.
Figure 13: An illustration of Retrieve Error: The retrieve system incorrectly identifies and returns an inappropriate tool.
Figure 14: An illustration of Tool Error: The LLM misapplication with improper tool parameters.
Figure 15: An illustration of Other Error: The agent’s inadequate response to tool timeout.

### v2

@@ E.2. Annotation Principles
Appendix F Details of LiveMCPTool Collection
F.1. Toolset Statistics
The statistics of LiveMCPTool’s tools and servers are illustrated in Figures 11-11. The collected toolset is categorized into eight distinct classes:
  1. (1)
@@ F.1. Toolset Statistics
F.2. Tool Verification for LiveMCPTool.
To address concerns regarding the dynamism of the LiveMCPTool, we conducted a systematic verification of all collected tool implementations. For tools that require online information (e.g., weather, news, or stock data), we confirmed that they indeed retrieve real data from online data sources rather than relying on simulated responses or pre-cached static datasets. This validation ensures that the server behaviors used in our benchmark faithfully reflect dynamic, real-world conditions, thereby supporting our claim of genuine runtime dynamism.
Figure 12. An illustrative case of evaluation failure in LiveMCPEval. The evaluator model erroneously concluded the task was successful based solely on the agent’s file creation action, while failing to recognize that the agent did not actually acquire the required information.
Appendix G Implementation Details
@@
  • OpenRouter: GPT-5, GPT-5-Mini, GPT-4.1, GPT-4.1-Mini, DeepSeek-R1-0528, DeepSeek-V3-0324, Qwen3-235B-A22B and Qwen3-32B.
@@
To address suboptimal greedy decoding in certain reasoning models, we implemented a uniform temperature parameter of 0.7 across all experiments. This configuration introduces controlled stochasticity while maintaining result reliability for long-horizon tasks, as we observed that sporadic randomness has a negligible cumulative impact on aggregate performance. We compute the total expenditure based on the model prices provided by OpenRouter as of January 31, 2026, accounting for both input and output tokens.
Appendix H Evaluation Analysis
H.1. Evaluation Metrics and Judge Model
In LiveMCPBench, we primarily adopt the single-run success rate as the main comparison metric. However, depending on different evaluation needs, alternative metrics may be more appropriate. For example, the pass@k metric (success in at least one out of $k$ independent samples) can be used to assess an agent’s ability to discover diverse solution paths. Conversely, the passˆk metric (success in all $k$ independent samples) evaluates the agent’s consistency and reliability.
For the judge model used in LiveMCPEval, we recommend employing DeepSeek-V3-0324 to obtain results comparable to those reported in this paper. Nevertheless, other models with higher human agreement or more robust evaluation frameworks may also be used, particularly by leveraging our annotated trajectories based on the Claude family.
Figure 13. Comparison of key points in DeepSeek-V3 and human annotations: Similar content despite different ordering.
H.2. Tool-Use Hallucination
Tool-use hallucination is a major source of evaluation failure. Because the evaluator judges task completion solely from the agent’s trajectory, hallucinated tool calls or fabricated tool outputs can mislead the evaluator. Therefore, a careful analysis and practical mitigation strategy are necessary.
Conventional rule-based, post-hoc verification is difficult to scale in a large and dynamic MCP environment. Different tool combinations (e.g., multiple search engines in the MCP ecosystem) can achieve equivalent functionality, making it hard to encode capability-equivalence checks with fixed rules. Likewise, answer–based verification is brittle under dynamic tasks such as daily news summarization. As a result, reliable evaluation protocols that are robust to hallucinations are challenging to deploy in real MCP settings.
A promising alternative is to have the evaluator decompose the task into explicit completion criteria and then locate supporting evidence directly from the raw tool outputs. Tool-use hallucinations most often arise when the agent invents intermediate information across multi-step reasoning, and the evaluator fails to notice the fabrication. Evidence-grounded, criterion-wise evaluation can substantially reduce this risk. This can be viewed as an extension of our key-points–based evaluation: each key point is verified independently against the tool sequence.
To quantify the prevalence of tool-use hallucination, we analyzed 814 tool invocations produced by Claude-Sonnet-4. Only 9.00% involved hallucination, indicating that such errors are relatively infrequent. To assess their impact on evaluation, we measured disagreement between the DeepSeek-V3 evaluator and human judgments due to tool-use hallucination across Claude-family models. The inconsistency rate was only 1.6%. Among 11 observed hallucination-induced errors, 8 were detected by the evaluator, yielding a 72.7% detection rate. Thus, the current protocol already catches most such failures, and the remaining undetected cases place a very small upper bound on evaluation.
Finally, we tested whether tool-use hallucinations are reliably detected across evaluators. We compared the distribution of evaluation errors when using Qwen2.5-72B as the evaluator against the default DeepSeek-V3. If detection were unstable, the error-type distribution would shift substantially. Instead, the distribution remained similar: Completeness Error (33.3%), Hallucinated Completion (46.2%), Granularity Mismatch (15.4%), and Others (5.1%). This close match suggests that hallucination-related failures are captured consistently rather than introducing evaluator-dependent bias. Therefore, tool-use hallucinations are unlikely to materially distort our evaluation results.
H.3. Case Study: Error Evaluation Examples
@@
H.4. Human Key Points
What constitutes a key point. A key point is an actionable, minimal step that contributes to solving the task using the MCP toolset. Each key point must (i) be atomic (one tool-mediated action per point), (ii) be anchored to a concrete tool or observable signal, (iii) be auditable from logs (inputs/outputs). Proposers write key points while completing the task, following the principles in Section E.
Normalization and consolidation. After collection, verifiers normalize key points into a canonical set by enforcing: (1) imperative voice; (2) one action per point; (3) explicit dependencies (e.g., “after obtaining X from Tool-A, call Tool-B with Y”); (4) de-duplication of paraphrases and overlapping sub-steps. When multiple proposers describe similar operations with different granularity, verifiers keep the coarsest level that remains auditable and does not obscure required tool interactions.
H.5. Case Study: Human and LLM Key Points Examples
@@
Appendix I Case Study: Error Examples
Figures 14-17 present concrete examples of four distinct error types: Query Error (13.33%), Retrieve Error (50.00%), Tool Error (18.33%), and Other Error (18.33%).
@@ Appendix I Case Study: Error Examples
Figure 14. An illustration of Query Error: Discrepancy between the agent-generated query and the task’s required competencies.
Figure 15. An illustration of Retrieve Error: The retrieve system incorrectly identifies and returns an inappropriate tool.
Figure 16. An illustration of Tool Error: The LLM misapplication with improper tool parameters.
Figure 17. An illustration of Other Error: The agent’s inadequate response to tool timeout.
Appendix J Prompts